# Keep It Real: Challenges in Attacking Compression-Based Adversarial Purification

### v1

@@ G.4 Iterative Defenses
Table 12: Accuracy (%) for iterative JPEG defenses as in Räber et al. [2025]. Performing the attack (PGD with epsilon 8/255) on a defense with just one iteration defeats the iterative version. The robust accuracy seems connected to the number of iterations in the attack, not the defense.
It was claimed in prior work that applying compression iteratively strengthens adversarial defenses [Räber et al., 2025]. However, we show that this effect is primarily due to gradient masking rather than true robustness (cf. Table˜12 in Section˜G.4). When attacking an iterative defense by approximating gradients using fewer defense iterations, the gradients remain informative, and accuracy can be reduced to levels comparable to using a single defense iteration.

### v3

@@ G.4 Iterative Defenses
Table 12: Accuracy (%) for iterative JPEG defenses as in [2]. Performing the attack (PGD with epsilon 8/255) on a defense with just one iteration defeats the iterative version. The robust accuracy seems connected to the number of iterations in the attack, not the defense.
It was claimed in prior work that applying compression iteratively strengthens adversarial defenses [2]. However, we show that this effect is primarily due to gradient masking rather than true robustness (cf. Table˜12 in Section˜G.4). When attacking an iterative defense by approximating gradients using fewer defense iterations, the gradients remain informative, and accuracy can be reduced to levels comparable to using a single defense iteration.